# Data-free Asymptotics-Informed Operator Networks for Singularly Perturbed PDEs

### v1

@@ 3.5 Comparison with the original FEONet
The original FEONet is limited by its reliance on $P_{1}$ and $P_{2}$ basis functions, which are insufficient for resolving stiff behaviors. As a result, it offers no significant performance advantage over the standard FEM (see the middle plot in Figure 11). To address this, we proposed eFEONet, which enriches the basis functions with exponential (exp) and error (erf) functions derived from the asymptotic analysis (see Appendix A for details), ensuring they align with the mathematical structure of the problem’s stiff behavior. This enhancement enables eFEONet to outperform the original FEONet, yielding substantially reduced errors across both boundary and interior layers, as shown in Table 6.
Figure 11: Visualization of the true solution and solution plots from FEONet and eFEONet of the boundary layer problem when $\varepsilon=1e-05$
@@ Acknowledgments
Y. Hong was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1A2C1093579) and by the Korea government(MSIT) (RS-2023-00219980). S. Ko was supported by National Research Foundation of Korea Grant funded by the Korean Government (RS-2023-00212227). Jae Yong Lee was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)).
@@ Appendix A Derivation of Corrector Basis Functions
where $0<\varepsilon\ll 1$, and $\boldsymbol{b}=\boldsymbol{b}(\mathbf{x})$, $c=c(\mathbf{x})$ and $f=f(\mathbf{x})$ are given smooth functions defined over the domain $D$. This formulation represents a general convection-diffusion-reaction equation with singular perturbation. For this problem, we considered both 1D and 2D settings, addressing critical challenges such as boundary layers and interior layers that arise due to the small parameter $\varepsilon>0$. From this point onward, our analysis follows the singular perturbation analysis stated in (BLA_3). The theoretical foundations and techniques presented here are based on this approach, providing a rigorous framework for handling boundary and interior layers in singularly perturbed problems. For further details on related studies and extensions, we refer the reader to (BLA_3).
(Boundary layer case) While our ultimate goal is to solve the above problem in 2D, we first simplify the analysis and explanation by considering a one-dimensional paradigm problem. The 1D problem is defined as
@@
(Interior layer case): For convection–diffusion equations with an interior layer, we consider the problem
@@
where $b(x)$ satisfies $b<0$ for $x<0$, $b(0)=0$, $b>0$ for $x>0$, and $b^{\prime}(x)>0$. The turning point at $x=0$ introduces an interior layer due to the change in sign of $b(x)$, where characteristics collide. For the formal limit problem, setting $\varepsilon=0$ leads to:
$b(x)u_{0}^{\prime}=f,$
but this may not be well-defined at $x=0$ since $b(0)=0$. Therefore, we split the solution into left and right parts, $u_{0}^{l}$ and $u_{0}^{r}$, corresponding to $x<0$ and $x>0$, respectively
$b(x)(u_{0}^{l})^{\prime}=f\quad\text{for }x<0\quad{\rm{and}}\quad-b(x)(u_{0}^{r})^{\prime}=f\quad\text{for }x>0.$
@@
where erf denotes an error function. This serves as a corrector for the interior layer.

### v2

@@ 3.5 Comparison with the original FEONet
The original FEONet is limited by its reliance on piecewise polynomial basis functions, which are insufficient for resolving stiff behaviors. As a result, it offers no significant performance advantage over the standard FEM (see the middle plot in Figure 11). To address this, we proposed eFEONet, which enriches the basis functions with exponential and error functions derived from the asymptotic analysis (see Appendix A for details), ensuring they align with the mathematical structure of the problem’s stiff behavior. This enhancement enables eFEONet to outperform the original FEONet, yielding substantially reduced errors across both boundary and interior layers, as shown in Table 6.
Figure 11: Visualization of the true solution and solution plots from FEONet and eFEONet of the boundary layer problem when $\varepsilon=10^{-5}$.
@@ Acknowledgments
Y. Hong was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1A2C1093579) and by the Korea government(MSIT) (RS-2023-00219980). S. Ko was supported by National Research Foundation of Korea Grant funded by the Korean Government (RS-2023-00212227). Jae Yong Lee was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)), and by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-02303239).
@@ Appendix A Derivation of Corrector Basis Functions
where $0<\varepsilon\ll 1$, and $\boldsymbol{b}=\boldsymbol{b}(\mathbf{x})$, $c=c(\mathbf{x})$ and $f=f(\mathbf{x})$ are given smooth functions defined over the domain $D$. This formulation represents a general convection-diffusion-reaction equation with singular perturbation. For this problem, we considered both 1D and 2D settings, addressing critical challenges such as boundary layers and interior layers that arise due to the small parameter $\varepsilon>0$. From this point onward, our analysis follows the singular perturbation analysis stated in (Gie et al., 2018). The theoretical foundations and techniques presented here are based on this approach, providing a rigorous framework for handling boundary and interior layers in singularly perturbed problems. For further details on related studies and extensions, we refer the reader to (Gie et al., 2018).
Boundary layer case. We simplify the analysis and explanation by considering a one-dimensional paradigm problem. The 1D problem is defined as
@@
Interior layer case. For convection-diffusion equations with an interior layer, we consider the problem
@@
where $b(x)$ satisfies $b<0$ for $x<0$, $b(0)=0$, $b>0$ for $x>0$, and $b^{\prime}(x)>0$. The turning point at $x=0$ introduces an interior layer due to the change in sign of $b(x)$, where characteristics collide. For the formal limit problem, setting $\varepsilon=0$ leads to
$b(x)u_{0}^{\prime}=f.$
However, it may not be well-defined at $x=0$ since $b(0)=0$. Therefore, we split the solution into left and right parts, $u_{0}^{l}$ and $u_{0}^{r}$, corresponding to $x<0$ and $x>0$, respectively
$b(x)(u_{0}^{l})^{\prime}=f\quad\text{for }x<0,\quad{\rm{and}}\quad-b(x)(u_{0}^{r})^{\prime}=f\quad\text{for }x>0.$
@@
where erf denotes a Gaussian error function. This serves as a corrector for the interior layer.